OpenStudy (anonymous):

Which of the following has the greatest mass, in grams?1 atom of gold (Au) 1.0 mol mercury (Hg) atoms 2.0 mole of chlorine (Cl) atoms 25 atoms of mercury (Hg)

OpenStudy (stephstar101):

1 atom Au - 196.96655g 1 mol Hg - 200.59g*(6.02*10^23 particles)= 1.208*10^26g 2 mol Cl - 35.453g*2*(6.02*10^23 particles)= 4.269*10^25g 25 atoms Hg - 25*200.59g= 5014.75g So the largest mass in g would be 1 mol Hg with a mass of 1.208*10^26g

OpenStudy (anonymous):

thank you so much!